OpenStudy (anonymous):

Find the distance between the points 2,-3 and 5,-4

OpenStudy (texaschic101):

d = sq rt (x2 - x1)^2 + (y2 - y1)^2 (2,-3) x1 = 2 and y1 = -3 (5,-4) x2 = 5 and y2 = -4 now we sub d = sq rt (5 - 2)^2 + (-4 - (-3)^2 d = sq rt 3^2 + (-1^2) d = sq rt (9 + 1) d = sq rt 10 d = 3.16